OpenStudy (anonymous):

Simplify. Assume that no denominator is equal to zero. r^6n^-7/r^4n^2

OpenStudy (cggurumanjunath):

a^b/a^c=a^(b-c)

OpenStudy (cggurumanjunath):

use the above rule of exponent.

OpenStudy (cggurumanjunath):

r^6/r^4 becomes r^(6-4)

OpenStudy (cggurumanjunath):

now do that for n similarly.

OpenStudy (cggurumanjunath):

n^(-7)/n^2=n^(-7-2)=n^(-9)

OpenStudy (cggurumanjunath):

so the final answer is the product of r and n; r^2*n^(-9)

OpenStudy (cggurumanjunath):

a^-x=1/a^(x)

OpenStudy (cggurumanjunath):

use the above rule

OpenStudy (cggurumanjunath):

r^2/n^9 is ur final answer.